OpenStudy (marmar10):

During the combustion of 5.000 g of octane, C8H18, 239.5 kcal is released. 2C8H18(l)+25O2(g)→16CO2(g)+18H2O(g) How much energy is released by the combustion of 1.010 mol of C8H18? How many grams of octane must be burned to release 427.3 kcal ? How many kilocalories are released by the combustion of 19.0 g of C8H18 ?

OpenStudy (marmar10):

having trouble with these :/

OpenStudy (marmar10):

@Photon336

OpenStudy (matt101):

Based on the information provided in the question, first calculate the enthalpy in kJ/mol. This value gives you a way to answer all the questions!

OpenStudy (photon336):

@Cuanchi

OpenStudy (cuanchi):

I think it is not necessary to convert to the kcal to kJ/mol, you can just use conversion factors. \[ \frac{5.000 g octane }{239.5 kcal } =\frac{ 239.5 kcal }{ 5.000 g octane } \] a) calculate the molecular mass of octane C8H18= 12x 8 + 18 = MM g octane 1.010 mol of C8H18 x (MM g octane/ 1 mol) x (239.5 kcal/ 5.000 g octane) = _____ kcal b) 427.3 kcal x (5.000 g octane/ 239.5 kcal) = ________ g of octane c) 19.0 g of C8H18 x (239.5 kcal/ 5.000 g octane) = _______kcal

OpenStudy (matt101):

Sorry, misread the kcal as kJ. You're right, no conversion necessary!